OpenStudy (anonymous):

How do you change a phasor from polar form to rectangular form? How does 3.7<36 degrees (< is the angle symbol) convert to 3+j2.16? I understand everything else but this is bothering me.

OpenStudy (anonymous):

yes. 3.7<36 is 2.9933+j2.1748. It can be converted by the following steps. (Assume < is the angle sign everywhere it occurs) 1. Polar form is \[r < \theta \] In your problem, \[r=3.7, \theta=36^{o}\] 2. Convert to rectangular using the formula \[x=r \cos \theta, y=r \sin \theta \] 3. Obtain the answer as \[r <\theta =x +jy\]